OpenStudy (ineedhelp101):

i need help with an equation for pre-cal. the equation is: x = 2 square root of (x-1)

OpenStudy (anonymous):

solve for x ?

OpenStudy (ineedhelp101):

yes

OpenStudy (anonymous):

square both sides: x^2=4(x-1) x^2=4x-4 x^2-4x+4=0 (x-2)^2=- x=2

OpenStudy (ineedhelp101):

thank you very much

OpenStudy (anonymous):

click "Good Answer" :)